OpenStudy (anonymous):

if f(x) = x + 4, find f(3). 1 7 −1 12

OpenStudy (anonymous):

need help n explain it to me

OpenStudy (anonymous):

If f(x) = x + 4, find f(3). y=3+4 y=7

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

ight

OpenStudy (anonymous):

yes he is right, it is like saying substitute x=3 into the equation.

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

f(x) = x + 4 f(3) = 3 + 4 = 7

OpenStudy (anonymous):

no problem.